OpenStudy (spicytaco1):

Simplify. 14 – 6.2 ∙ (–3 + –7) A. –78 B. –48 C. 76 D. 78

OpenStudy (anonymous):

14 – 6.2 ∙ (–3 + –7) =14 - 6.2 times (-10) = 14 - (-62) =14+62 =76

OpenStudy (anonymous):

Of course C!